OpenStudy (anonymous):

Which of these is the composite function f(g(x))? f(x) = x2 – 5 g(x) = 9x – 2

OpenStudy (anonymous):

f(g(x)) = 81x^2 – 18x – 9 f(g(x)) = 81x^2 – 36x – 1 f(g(x)) = 9x2 – 9 f(g(x)) = 9x^2 – 1